# BiDoRA: Bi-level Optimization-Based Weight-Decomposed Low-Rank Adaptation

### v1

@@ 5.5 Ablation Studies
In this section, we perform ablation studies to investigate the effectiveness of individual modules or strategies in BiDoRA. We fine-tune a RoBERTa-base model on the GLUE benchmark under different ablation settings, and the results are shown in Table 5.
Retraining          We test the model directly obtained from the search phase to evaluate the effectiveness of further retraining the incremental direction component. The results show that BiDoRA outperforms BiDoRA (w/o retraining) on average, highlighting the necessity of retraining.
Bi-level Optimization          We set $\xi$ to zero in Algorithm 1 to assess the effectiveness of the bi-level optimization framework. This ablation setting can be interpreted as an alternative learning method where two optimization steps are carried out alternately on two different splits of the training dataset. Notably, in the alternative learning method, the updating of each component is unaware of each other, making the training less stable. In contrast, the hyper-gradient used in bi-level optimization avoids this issue by connecting the two levels in a certain way. The results show that BiDoRA outperforms BiDoRA ($\xi=0$) on average, demonstrating the efficacy of the bi-level optimization strategy.
Orthogonal Regularization          We examine the effectiveness of the orthogonality constraint in Eq. (3) by setting $\gamma$ to zero. Results show that BiDoRA outperforms BiDoRA (w/o cst.) on average, indicating the effectiveness of applying the orthogonality regularizer to alleviate overfitting.
Table 5: Ablation studies. We evaluate the performance of BiDoRA without retraining (w/o retraining), without bi-level optimization ($\xi=0$) and without orthogonal regularization (w/o cst.).
Method MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg
BiDoRA (w/o retraining) $87.0$ $94.2$ $89.0$ $57.3$ $92.4$ $90.6$ $71.6$ $90.0$ $84.0$
BiDoRA ($\xi=0$) $86.9$ $94.2$ $89.0$ $59.4$ $90.8$ 91.2 $75.9$ $90.0$ 84.7
BiDoRA (w/o cst.) $87.0$ 94.4 $88.6$ 61.3 92.7 $90.2$ $76.0$ 90.1 $85.0$
BiDoRA 87.1 94.4 89.4 61.3 92.7 $90.6$ 76.1 90.1 85.2
@@ 5.6 Weight Decomposition Analysis
One important motivation of DoRA is to bridge the inherent differences between LoRA and FT. Similar to DoRA, we conduct a weight decomposition analysis on the correlation between the change of magnitudes and that of directions (detailed in Appendix E) for BiDoRA and baseline methods by fine-tuning a GPT2-medium model on the E2E dataset. As shown in Figure 3, FT, DoRA, and BiDoRA all exhibit negative correlation values, while LoRA shows a positive correlation, consistent with the findings in Liu et al. (2024). Notably, BiDoRA achieves a negative correlation of $-8.042$, closer to FT than DoRA’s $-1.784$. This improvement is attributed to the decoupled training process of the two layers, which allows for a higher learning capacity compared to DoRA.
5.7 Computation Costs
Table 6: Average training time cost on the MNLI, QQP, and SST-2 datasets.
Method LoRA DoRA BiDoRA
Cost ×$\times 1$$\times 1.30$$\times 3.92$
Since BiDoRA has the same architecture as DoRA, our method only requires two extra forward and backward passes of the lower level for the hypergradient calculation of the upper level, as shown in Eq. 8. In theory, this would make our method roughly threefold computationally costly. A similar analysis holds for the memory consumption analysis. Empirically, table 6 shows the average training cost of BiDoRA and two baseline methods on the MNLI, QQP, and SST-2 datasets from the GLUE benchmark, being consistent with the theoretical analysis. We normalize the cost of LoRA to $1$ for reference. Given BiDoRA’s superior performance across various tasks, the increase in computational costs is acceptable in most cases, underscoring its practicality.
6 Conclusion
@@
We propose BiDoRA, a novel bi-level optimization framework for parameter-efficient fine-tuning of large-scale pre-trained models. By conducting weight decomposition following the DoRA approach, our method trains the two components separately in two interconnected optimization levels using different sub-datasets. In this way, BiDoRA not only decouples the learning process of the two components, resulting in a learning pattern closer to FT, but also effectively alleviates overfitting. Empirical studies on various NLP tasks demonstrate that BiDoRA outperforms DoRA and other baselines, highlighting the effectiveness of our method.
@@ Appendix A Datasets and Models
A.1 Natural Language Understanding
The GLUE Benchmark comprises a diverse array of tasks that are widely employed for evaluation in natural language understanding. It encompasses two single-sentence classification tasks, three tasks assessing similarity and paraphrasing, and four tasks focusing on natural language inference. Specifically, it includes MNLI (MultiNLI, Williams et al. (2017)), SST-2 (Stanford Sentiment Treebank, Socher et al. (2013)), MRPC (Microsoft Research Paraphrase Corpus, Dolan & Brockett (2005)), CoLA (Corpus of Linguistic Acceptability, Warstadt et al. (2019)), QNLI (Question NLI, Rajpurkar et al. (2018)), QQP (Quora Question Pairs), RTE (Recognizing Textual Entailment), and STS-B (Semantic Textual Similarity Benchmark, Cer et al. (2017)). We summarize the statistical data for all datasets within the GLUE Benchmark in Table 7.
The Reuters-21578 (Padmanabhan et al., 2016) dataset is one of the most widely used data collections for text categorization research. It was collected from the Reuters financial newswire service in 1987 and is used for text classification and natural language processing tasks. Three splits are available: ModApte, ModHayes, and ModLewis. These documents cover various topics, such as politics, economics, and sports. We summarize the statistical data for all text classification tasks used in our experiments in Table 8.
Table 7: The statistical data for all datasets within the GLUE Benchmark.
Dataset Metrics Train Dev Test Label Task
MNLI Accuracy 393k 20k 20k 3 NLI
SST-2 Accuracy 67k 872 1.8k 2 Sentiment
MRPC Accuracy 3.7k 408 1.7k 2 Paraphrase
CoLA Matthews Corr 8.5k 1k 1k 2 Acceptability
QNLI Accuracy 108k 5.7k 5.7k 2 QA/NLI
QQP Accuracy 364k 40k 391k 2 Paraphrase
RTE Accuracy 2.5k 276 3k 2 NLI
STS-B Pearson Corr 7.0k 1.5k 1.4k 1 Similarity
Table 8: The statistical data for the Reuters-21578 dataset.
Dataset Metrics Train Test
ModApte F1 8.8k 3k
ModHayes F1 18k 0.7k
ModLewis F1 12k 5.5k
A.2 Natural Language Generation
In our experiments on natural language generation, we use the E2E (Novikova et al., 2017) dataset, which was initially introduced as a dataset for training end-to-end, data-driven natural language generation systems. Multiple references can be associated with each source table used as input. Each sample input $(x,y)$ consists of a series of slot-value pairs accompanied by an associated natural language reference text. The E2E dataset comprises approximately 42,000 training examples, 4,600 validation examples, and 4,600 test examples from the restaurant domain.
We utilize the following five evaluation metrics: BLEU (Papineni et al., 2002), NIST (Lin & Och, 2004), METEOR (Banerjee & Lavie, 2005), ROUGE-L (Lin, 2004), and CIDEr (Vedantam et al., 2015). We summarize its statistical data in Table LABEL:tab:datasetnlg.
Table 9: The statistical data for E2E dataset.
Dataset Metrics Train Validation
E2E BLEU,NIST,MET,ROUGE-L,CIDEr 42k 4.6k
A.3 Token Classification
BioNLP (Collier et al., 2004) is a Named Entity Recognition dataset that contains biological entities such as DNA, RNA, and protein. It is essentially a token classification task where we want to classify each entity in the sequence. CoNLL-2003 (Sang & De Meulder, 2003) focuses on language-independent named entity recognition. It concentrates on four types of named entities: persons, locations, organizations, and miscellaneous entities that do not belong to the previous three groups. We summarize the statistical data for all used token classification tasks in Table 10.
Table 10: The statistical data for token classification tasks
Dataset Metrics Train Validation Test
@@ Appendix B Experimental Settings
In this section, we provide detailed experimental settings. We maintain consistent configurations across experiments, including LoRA rank, LoRA $\alpha$, batch size, maximum sequence length, and optimizer, to ensure a fair comparison. The hyperparameter tuning for our method is straightforward and convenient.
@@ B.1 RoBERTa
We summarize the experimental settings for the GLUE benchmark in Table 11 and for the Reuters21578 dataset and token classification tasks in Table 12.
Table 11: The hyperparameters we used for RoBERTa on the GLUE benchmark.
Method Settings MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
@@
LoRA rank ranka = ranku = $4$
LoRA $\alpha$ $8$
RoBERTa-base Total batch size $32$
Global steps $20000$ $12000$ $25000$ $20000$ $15000$ $20000$ $15000$ $12000$
Lower learning rate 5e-5 1e-5 2e-5 5e-5 2e-5 5e-5 1e-5 1e-5
Upper learning rate 5e-5 1e-5 2e-5 5e-5 2e-5 5e-5 1e-5 1e-5
Lower weight decay $0.1$
Upper weight decay $0.1$ $0.1$ $0.1$ $0.1$ 00 $0.1$ $0.1$ $0.01$
Max Seq Length $512$
Regularization Coefficient 1e-5
RoBERTa-large Total batch size $32$
Global steps $50000$ $20000$ $30000$ $20000$ $60000$ $40000$ $15000$ $10000$
Lower learning rate 1e-5
Upper learning rate 1e-5
Lower weight decay $0.5$ $0.5$ 00 $0.2$ $0.5$ $0.5$ $0.5$ $0.5$
Upper weight decay $0.5$ $50.0$ $0.2$ $0.5$ $0.5$ $0.1$ $0.5$
Max Seq Length $128$
Regularization Coefficient 0 0 1e-5 1e-5 0 1e-5 0 1e-5
Table 12: The hyperparameters we used for RoBERTa on the Reuters21578 dataset, BioNLP dataset, and CoNLL2003 dataset.
Method Settings ModApte ModHayes ModLewis BioNLP CoNLL2003
Optimizer AdamW
Warmup Ratio $0.06$
Scheduler Linear
LoRA rank ranka = ranku = $4$
LoRA $\alpha$ $8$
RoBERTa-base Total batch size $32$
Global steps $20000$ $20000$ $20000$ $12000$ $12000$
Lower learning rate 3e-5 3e-5 3e-5 1e-5 2e-5
Upper learning rate 3e-5 3e-5 3e-5 1e-5 2e-5
Lower weight decay $0.1$ $0.1$ $0.1$ $0.1$ $0.2$
Upper weight decay $0.1$
Max Seq Length $512$
Regularization Coefficient 0 1e-5 0 1e-5 0
RoBERTa-large Total batch size $32$
Global steps $20000$ $20000$ $20000$ $12000$ $15000$
Lower learning rate 1e-5 1e-5 1e-5 2e-5 1e-5
Upper learning rate 1e-5 1e-5 1e-5 2e-5 1e-5
Lower weight decay $0.2$ $0.1$ $0.2$ $0.02$ $0.1$
Upper weight decay $0.1$ $0.1$ $0.1$ $0.02$ $0.1$
Max Seq Length $128$
Regularization Coefficient 0 1e-5 0 0 1e-5

### v2

@@ 5.4 Ablation Studies
Table 5: Ablation studies. We evaluate the performance of BiDoRA without retraining (w/o retraining), without BLO ($\xi=0$), without orthogonal regularization (w/o cst.), and with retraining magnitude.
Method MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg.
BiDoRA (retraining magnitude) 87.087.087.0 94.394.394.3 89.189.189.1 60.760.760.7 92.7 91.091.091.0 73.473.473.4 89.989.989.9 84.884.884.8
BiDoRA (w/o retraining) 87.087.087.0 94.294.294.2 89.089.089.0 57.357.357.3 92.492.492.4 90.690.690.6 71.671.671.6 90.090.090.0 84.084.084.0
BiDoRA ($\xi=0$) 86.986.986.9 94.294.294.2 89.089.089.0 59.459.459.4 90.890.890.8 91.2 75.975.975.9 90.090.090.0 84.7
BiDoRA (w/o cst.) 87.087.087.0 94.4 88.688.688.6 61.3 92.7 90.290.290.2 76.076.076.0 90.1 85.085.085.0
BiDoRA 87.1 94.4 89.4 61.3 92.7 90.690.690.6 76.1 90.1 85.2
In this section, we perform ablation studies to investigate the effectiveness of individual modules or strategies in BiDoRA. We fine-tune a RoBERTa-base model on the GLUE benchmark under different ablation settings, and the results are shown in Table˜5.
Retraining.
We test the model directly obtained from the search phase to evaluate the effectiveness of further retraining the direction component. The results show that BiDoRA outperforms BiDoRA (w/o retraining) on average, highlighting the necessity of retraining. Table˜5 also validates that retraining the direction component leads to superior performance than retraining the magnitude.
Bi-level optimization.
We set $\xi$ to zero in Algorithm˜1 to assess the effectiveness of the BLO framework. This ablation setting can be interpreted as an alternative learning method where two optimization steps are carried out alternately on two different splits of the training dataset. Notably, in the alternative learning method, the updating of each component is unaware of the others, making the training less stable. In contrast, the hyper-gradient used in BLO avoids this issue by connecting the two levels in a certain way. The results show that BiDoRA outperforms BiDoRA ($\xi=0$) on average, demonstrating the efficacy of the BLo strategy.
Orthogonal regularization.
We examine the effectiveness of the orthogonality constraint in Eq.˜3 by setting $\gamma$ to zero. Results show that BiDoRA outperforms BiDoRA (w/o cst.) on average, indicating the effectiveness of applying the orthogonality regularizer to alleviate overfitting.
@@ 5.5 Weight Decomposition Analysis
One important motivation of DoRA is to bridge the inherent differences between LoRA and FT. Similar to DoRA, we conduct a weight decomposition analysis on the correlation between the change of magnitudes and that of directions (detailed in Appendix˜D) for BiDoRA and baseline methods by fine-tuning a GPT2-medium model on the E2E dataset. As shown in Fig.˜3, FT, DoRA, and BiDoRA all exhibit negative correlation values, while LoRA shows a positive correlation, consistent with the findings in Liu et al. (2024a). Notably, BiDoRA achieves a negative correlation of 8.042-8.042- 8.042, closer to FT than DoRA’s 1.784-1.784- 1.784. This improvement is attributed to the decoupled training process of the two layers, which allows for a higher learning capacity compared to DoRA.
[…]
Figure 3: Magnitude and direction updates for (a) FT, (b) LoRA, (c) DoRA, and (d) BiDoRA of the query matrices across different layers and intermediate steps after fine-tuning the GPT2 model on the E2E dataset (Novikova et al., 2017), where $k$ denotes the correlation value. Different markers represent matrices from different training steps, with each color corresponding to a specific layer. $\Delta\mathbf{M}$ denotes the average change in weight vector magnitude, and $\Delta\mathbf{D}$ denotes the average change in direction, as formally defined in Appendix˜D.
5.6 Discussion
The advantage of BiDoRA is supported by both theoretical insights and empirical evidence, as detailed as follows.
Motivation.
Theoretically, Liu et al. (2024a) showed that LoRA’s training pattern tends to be coupled in terms of magnitude-direction correlation, which degrades learning capacity. Their solution was to introduce a reparameterization that decouples these components in the formulation. We build upon DoRA following their theory and further decouple magnitude and direction in terms of training dynamics. Specifically, the two components are trained in separate loops within a bilevel optimization framework, which is expected to improve performance in an intuition similar to DoRA.
Empirical evidences.
We performed a Wilcoxon signed-rank test to compare the performance of DoRA and BiDoRA. Specifically, we used the results from Table˜1. For each PEFT method, we collected 999 values (888 values from each dataset plus the average performance) from one base model. We concatenated the results from three base models (RoBERTa-base, RoBERTa-large, and DeBERTa-XXL) to obtain a list of 272727 values. A comparison of these 272727 values between DoRA and BiDoRA reveals that BiDoRA is significantly better than DoRA, with a p-value of $2.4\times 10^{-4}$. This result demonstrates that BiDoRA offers a non-marginal improvement over DoRA.
Additionally, the weight decomposition analysis, including (Fig.˜3 and Fig.˜4), indicates that BiDoRA achieves better decoupling of the components compared to DoRA. Evaluation metrics across various tasks demonstrate the superior performance of BiDoRA, confirming that our decoupled optimization loop leads to improved outcomes.
6 Conclusion and Future Works
@@
We propose BiDoRA, a novel bi-level optimization framework for PEFT of large-scale pre-trained models. By conducting weight decomposition following the DoRA approach, our method trains the two components separately in two interconnected optimization levels using different sub-datasets. In this way, BiDoRA not only decouples the learning process of the two components, resulting in a learning pattern closer to FT, but also effectively alleviates overfitting. Empirical studies on various NLP tasks demonstrate that BiDoRA outperforms DoRA and other baselines, highlighting the effectiveness of our method.
One potential limitation of BiDoRA is its training efficiency (see Appendix˜E) in terms of per-step cost, which could be reduced by using more advanced hyper-gradient estimators, such as SAMA (Choe et al., 2023a) or MixFlow-MG (Kemaev et al., 2025). Furthermore, while we have empirically shown that BiDoRA induces better decoupling between the magnitude and direction components (Fig.˜3 and Fig.˜4), a formal theoretical analysis of this property is currently lacking and serves for future work.
Acknowledgments
P.X. acknowledges funding support from NSF IIS2405974, NSF IIS2339216, NIH R35GM157217, and NIH R21GM154171.
@@ Appendix A Datasets and Models
Table 6: Summary of datasets used in the experiments
Task Group Dataset Metrics Train Dev / Val Test
Natural Language
Understanding
MNLI Accuracy 393k 20k 20k
SST-2 Accuracy 67k 872 1.8k
MRPC Accuracy 3.7k 408 1.7k
CoLA Matthews Corr 8.5k 1k 1k
QNLI Accuracy 108k 5.7k 5.7k
QQP Accuracy 364k 40k 391k
RTE Accuracy 2.5k 276 3k
STS-B Pearson Corr 7.0k 1.5k 1.4k
Text
Classification
ModApte F1 8.8k - 3k
ModHayes F1 18k - 0.7k
ModLewis F1 12k - 5.5k
Natural Language
Generation
E2E BLEU, NIST, MET, ROUGE-L, CIDEr 42k 4.6k -
Token
Classification
@@
Biomedical
Experiments
Thermostability
prediction
Accuracy, Precision, Recall, F1 3,695 - 924
BBP Accuracy, Precision, Recall, F1 200 - 38
MIC MSE 936 - 104
In this section, we present the datasets and models used in experiments, and summarize the statistical data in Table˜6.
A.1 Natural Language Understanding
The GLUE Benchmark (Wang et al., 2019) comprises a diverse array of tasks that are widely employed for evaluation in natural language understanding. It encompasses two single-sentence classification tasks, three tasks assessing similarity and paraphrasing, and four tasks focusing on natural language inference. Specifically, it includes MNLI (MultiNLI, Williams et al. (2018)), SST-2 (Stanford Sentiment Treebank, Socher et al. (2013)), MRPC (Microsoft Research Paraphrase Corpus, Dolan & Brockett (2005)), CoLA (Corpus of Linguistic Acceptability, Warstadt et al. (2019)), QNLI (Question NLI, Rajpurkar et al. (2018)), QQP (Quora Question Pairs, Wang et al. (2017)), RTE (Recognizing Textual Entailment, Dagan et al. (2005)), and STS-B (Semantic Textual Similarity Benchmark, Cer et al. (2017)). We summarize the statistical data for all datasets within the GLUE Benchmark in Table˜6. Following existing practices, the development set is used in GLUE as the test data since the actual test set is not publicly available. We report the overall (matched and mismatched) accuracy for MNLI, Matthew’s correlation for CoLA, Pearson correlation for STS-B, and accuracy for the other tasks.
The Reuters-21578 (Padmanabhan et al., 2016) dataset is one of the most widely used data collections for text categorization research. It was collected from the Reuters financial newswire service in 1987 and is used for text classification and natural language processing tasks. Three splits are available: ModApte, ModHayes, and ModLewis. These documents cover various topics, such as politics, economics, and sports. F1 score is used as the evaluation metric across all three splits.
A.2 Natural Language Generation
In our experiments on natural language generation, we use the E2E (Novikova et al., 2017) dataset, which was initially introduced as a dataset for training end-to-end, data-driven natural language generation systems. Multiple references can be associated with each source table used as input. Each sample input $(x,y)$ consists of a series of slot-value pairs accompanied by an associated natural language reference text. The E2E dataset comprises approximately 424242k training examples, 4,6004,6004 , 600 validation examples, and 4,6004,6004 , 600 test examples from the restaurant domain.
We utilize the following five evaluation metrics: BLEU (Papineni et al., 2002), NIST (Lin & Och, 2004), METEOR (Banerjee & Lavie, 2005), ROUGE-L (Lin, 2004), and CIDEr (Vedantam et al., 2015).
A.3 Token Classification
For token classification, we fine-tune the RoBERTa-base and RoBERTa-large models on the BioNLP dataset (Collier et al., 2004) and the CoNLL2003 dataset (Tjong Kim Sang, 2002). BioNLP (Collier et al., 2004) is a Named Entity Recognition dataset that contains biological entities such as DNA, RNA, and protein. It is essentially a token classification task where we want to classify each entity in the sequence. CoNLL-2003 (Tjong Kim Sang, 2002) focuses on language-independent named entity recognition. It concentrates on four types of named entities: persons, locations, organizations, and miscellaneous entities that do not belong to the previous three groups. Accuracy, precision, recall, and F1 score are used as evaluation metrics.
A.4 Biomedical Experiments
The ESM (Evolutionary Scale Modeling, Rives et al. (2021)) model is a transformer-based protein language model designed for protein sequence analysis, leveraging the transformer architecture to capture evolutionary patterns. We fine-tune the ESM model using the Protein Aligner checkpoint (Zhang et al., 2024a) on two classification tasks—thermostability prediction (Chen et al. (2023), 936 training samples) and blood-brain barrier peptide prediction (BBP, Dai et al. (2021), 200 training samples)—and one regression task, minimum inhibitory concentration prediction (MIC, Ledesma-Fernandez et al. (2023), 3,695 training samples). Notably, protein analysis datasets are typically much smaller than those in NLP, in which case the large pre-trained models are prone to overfitting, even when using PEFT methods. The trainable parameters (on the order of millions) are significantly overparameterized compared to the available samples (thousands or even hundreds), highlighting the need for our overfitting-resilient counterpart.
@@ Appendix B Experimental Settings
In this section, we provide detailed experimental settings. We maintain consistent configurations across experiments, including LoRA rank, LoRA $\alpha$, batch size, maximum sequence length, and optimizer, to ensure a fair comparison. For results other than Table˜2, we do not include the bias term in PEFT linear layers. The hyperparameter tuning for our method is straightforward and convenient.
@@ B.1 RoBERTa
We summarize the experimental settings for the GLUE benchmark (Table˜1) and for the Reuters21578 dataset and token classification (Table˜13) tasks in Table˜7.
Table 7: The hyperparameters used for RoBERTa on the GLUE benchmark (Wang et al., 2019), Reuters21578 dataset (Padmanabhan et al., 2016), BioNLP dataset (Collier et al., 2004), and CoNLL2003 dataset (Tjong Kim Sang, 2002).
Settings MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B ModApte ModHayes ModLewis BioNLP CoNLL2003
@@
LoRA rank rank = 444
LoRA $\alpha$ 888
RoBERTa-base
Total batch size 323232
Global steps 202020k 121212k 252525k 202020k 151515k 202020k 151515k 121212k 202020k 202020k 202020k 121212k 121212k
Lower learning rate 5e-5 1e-5 2e-5 5e-5 2e-5 5e-5 1e-5 1e-5 3e-5 3e-5 3e-5 1e-5 2e-5
Upper learning rate 5e-5 1e-5 2e-5 5e-5 2e-5 5e-5 1e-5 1e-5 3e-5 3e-5 3e-5 1e-5 2e-5
Lower weight decay 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.20.20.2
Upper weight decay 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0 0.10.10.1 0.10.10.1 0.010.010.01 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1
Max Seq Length 512512512
Regularization Coefficient 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 0 1e-5 0 1e-5 0
RoBERTa-large
Total batch size 323232
Global steps 505050k 202020k 303030k 202020k 606060k 404040k 151515k 101010k 202020k 202020k 202020k 121212k 151515k
Lower learning rate 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 2e-5 1e-5
Upper learning rate 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 2e-5 1e-5
Lower weight decay 0.50.50.5 0.50.50.5 0 0.20.20.2 0.50.50.5 0.50.50.5 0.50.50.5 0.50.50.5 0.20.20.2 0.10.10.1 0.20.20.2 0.020.020.02 0.10.10.1
Upper weight decay 0.50.50.5 0.050.050.05 0 0.20.20.2 0.50.50.5 0.50.50.5 0.10.10.1 0.50.50.5 0.10.10.1 0.10.10.1 0.10.10.1 0.020.020.02 0.10.10.1
Max Seq Length 128128128
Regularization Coefficient 0 0 1e-5 1e-5 0 1e-5 0 1e-5 0 1e-5 0 0 1e-5